# End-to-End Crystal Structure Prediction from Powder X-Ray Diffraction

### v2

@@ Abstract
Crystal structure prediction (CSP) has made significant progress, but most methods focus on unconditional generations of inorganic crystal with limited atoms in the unit cell. This study introduces XtalNet, the first equivariant deep generative model for end-to-end CSP from Powder X-ray Diffraction (PXRD). Unlike previous methods that rely solely on composition, XtalNet leverages PXRD as an additional condition, eliminating ambiguity and enabling the generation of complex organic structures with up to 400 atoms in the unit cell. XtalNet comprises two modules: a Contrastive PXRD-Crystal Pretraining (CPCP) module that aligns PXRD space with crystal structure space, and a Conditional Crystal Structure Generation (CCSG) module that generates candidate crystal structures conditioned on PXRD patterns. Evaluation on two MOF datasets (hMOF-100 and hMOF-400) demonstrates XtalNet’s effectiveness. XtalNet achieves a top-10 Match Rate of 90.2% and 79% for hMOF-100 and hMOF-400 datasets in conditional crystal structure prediction task, respectively. XtalNet represents a significant advance in CSP, enabling the prediction of complex structures from PXRD data without the need for external databases or manual intervention. It has the potential to revolutionize PXRD analysis. It enables the direct prediction of crystal structures from experimental measurements, eliminating the need for manual intervention and external databases. This opens up new possibilities for automated crystal structure determination and the accelerated discovery of novel materials.
1 Main
Crystal structure plays a critical role in many scientific realms such as physics, chemistry and material science, as various physical and chemical properties of materials are derived from the crystal structure [1]. Crystal structure prediction (CSP), which aims to obtain the three-dimensional structure of crystals based on their composition [2], has made significant progress [3, 4, 5, 6]. These methods are carefully designed to account for the unique periodicity and equivariance challenges in CSP and some of them[3, 4, 5] utilized diffusion model [7, 8] to achieve remarkable structure generation results. However, these methods all conduct experiments on inorganic crystal structure dataset, such as Perov-5 [9, 10], MP-20 [11] and ICSD [12], with limited number of atoms in unit cell (usually less than 50). The efficacy of these methods in predicting more complex and larger organic structures or under stringent conditions has yet to be validated. As the number of atoms in unit cell increases, one chemical composition may exist multiple stable crystal structures resulting ambiguity for traditional CSP problem. On the other hand, in practice, crystal structure of a specific material is determined through experimental characterization, thereby establishing a one-to-one correspondence between the structure and experimental characterization. Therefore, integrating experimental characterization with crystal structure prediction is a more reasonable solution for larger crystal system.
Among the various experimental characterization techniques, such as Powder X-ray Diffraction (PXRD) and scanning transmission electron microscopy (STEM), PXRD is most widely used for its cost-effectiveness and ability to analyze various properties and determine crystal structures. However, the analysis of PXRD data often requires extensive human manual intervention, resulting numerous works try to automate PXRD analysis [13, 14, 15, 16, 17, 18]. While all these efforts only achieved some success at coarse-grained levels, such as phase identification and space group prediction, the more difficult and important task of fine-grained crystal structure prediction from PXRD remains unaddressed.
@@ 1 Main
In the practice of PXRD analysis, determining the crystal structure of experimental materials is one of the most important and challenging steps. Existing PXRD-based crystal structure determination methods first comparing experimental PXRD patterns of unknown materials against databases like the Inorganic Crystal Structure Database (ICSD) [19] helps identify structural analogue, which serves as the starting point for subsequent process. Then Rietveld refinement [20] is employed for refining rough analogous structures incrementally to achieve greater precision. However, the incompleteness of the database and the coarse nature of the matching results complicate the structure determination process, while the Rietveld refinement step often requires extensive manual intervention from experienced scientist. Therefore, a method that can directly predict crystal structures from PXRD data without relying on external databases has the potential to significantly reduce the manual labor required by chemists and represent a major step forward in automating PXRD analysis.
Despite the relatively underdeveloped state of crystal structure prediction from PXRD, the counterparts have been successfully implemented in the text-to-image generation domain, with prominent examples such as Stable Diffusion [21]. These models employ text embeddings generated by pretrained CLIP [22] text encoder to guide the diffusion process in generating images conditioned on textual descriptions. In this case, text guidance can be analogous to PXRD, while image generation can be analogous to crystal structure prediction. Drawing inspiration from the success models, we have explored the potential of integrating a contrastive learning approach and diffusion-based model into our conditional crystal structure prediction methodology. The contrastive PXRD-crystal pre-training strategy is applied to guide the generation of crystal structures based on PXRD pattern.
Therefore, we propose XtalNet, the first equivariant deep generative model for end-to-end crystal structure prediction from PXRD, strives to extend the capabilities of deep learning in generating crystal structures based on PXRD pattern, encompassing more complex structure and specific conditions. Unlike previous CSP methods that obtain crystal structures solely based on its composition, XtalNet adds PXRD as supplemental condition to make the CSP problem become a one-to-one mapping without ambiguity. XtalNet is an end-to-end method without human intervention or external database dependencies, which is departure from traditional PXRD analysis approaches. Our model is capable of handling organic crystal system, such as mental-orgnic framework (MOF) which is important for gas separation[23], with large number of atoms in unit cell (up to 400). This is made possible through a contrastive PXRD-crystal pre-training (CPCP) module and a conditional crystal structure generation (CCSG) module. The CPCP module utilize contrastive learning pre-training to align PXRD space with crystal structure space. Furthermore, by utilizing CCSG modules, multiple candidate crystal structures can be generated conditioned on PXRD pattern. Subsequently, these candidate structures are scored and ranked using the CPCP module, thus accomplishing the ranked structure prediction task with a top-10 match rate of 90.2% in hMOF-100 dataset and 79% in hMOF-400 dataset.
@@ 2.1 Overview of XtalNet
Figure 1: Over view of XtalNet. a, the framework of the Contrastive PXRD-Crystal Pretraining (CPCP) module. The CPCP module accepts PXRD patterns and crystal structures as inputs and outputs similarity scores between them. A transformer-based PXRD feature extractor processes the PXRD pattern data, while an equivariant GNN extracts features from the crystal structure. The dot product of these two features yields the similarity score. b, the framework of the Conditional Crystal Structure Generation (CPSG) module. The CCSG module employs the PXRD pattern as a condition to generate crystal structures. The PXRD feature extractor, initialized from the CPCP module pretraining and frozen, uses the composition of the crystal to initialize the atom positions and lattice matrix. Subsequently, the denoising network, i.e., the crystal structure network, takes the current crystal structure alongside the PXRD feature and time step as inputs to generate the update step. This process is iteratively repeated in reverse process. The diffusion process aims to add Gaussian noise into crystal structure resulting training denosing network. c, the PXRD data generation workflow. PXRD data can be obtained in two ways: either by simulating the PXRD data from a given crystal structure using GSAS[24] software or by conducting an actual PXRD experiment with an XRD instrument.
@@
XtalNet is designed for predicting crystal structure from PXRD, while we regard it as a conditional generation task. We generate the corresponding crystal structure based on given PXRD condition. To achieve this objective, we design two modules as depicted in Fig. 1a and b: the Contrastive PXRD-Crystal Pretraining (CPCP) module, and the Conditional Crystal Structure Generation (CCSG) module based on conditional diffusion. The CPCP module is primarily employed for aligning crystal space with PXRD space and can be utilized for database retrieval tasks, which aim to derive the most similar crystal structure based on a known PXRD pattern. Conversely, the CCSG module is crafted for recovering the crystal structure according to a given PXRD pattern using a pre-trained PXRD feature extractor, which is derived from the CPCP module.
The CPCP module adopts a design akin to CLIP[22]. A transformer-based PXRD feature extractor is employed to extract PXRD features. Subsequently, a crystal structure feature network based on Equivariant Graph Neural Networks is used to extract crystal structure features. Similarity scores are obtained by the cosine similarity between PXRD features and crystal structure features. Matching pairs of PXRD patterns and corresponding crystal structures are considered as positive pairs, whereas non-matching pairs are regarded as negative pairs. The infoNCE loss function is then adopted for contrastive learning training.
The CCSG module primarily utilizes a diffusion-based framework, with the PXRD feature serving as a crucial condition for generation. The noised crystal structure, the PXRD feature, and time embedding are taken by the crystal structure network to denoise fractional coordinates and lattice. The PXRD feature extractor is initialized by the CPCP module pretraining, which is vital for generation performance.
@@
To accomplishing XtalNet training, we must obtain pairs of crystal structure and PXRD; the entire workflow is illustrated in Fig. 1c. While crystal structure is prevalent in various databases, PXRD patterns are not ordinarily included in standard databases. Consequently, we simulate PXRD patterns from known crystal structures through GSAS [24] software as our training data. Simultaneously, we also employ X-ray diffraction instruments to collect experimental PXRD patterns to test XtalNet performance in practice.
@@ 2.2 Dataset Preparing and Evaluation Metrics
For the database retrieval task, we employ the top-$k$ hit rate as the evaluation metric, which measures the frequency at which the desired crystal structures are discovered. Following previous works[3, 4], we assess the ranked structure generation task using Match Rate and RMSE. Specifically, the Match Rate denotes the proportion of matched generated structures relative to all ground truth, while the StructureMatcher class in pymatgen [27] with thresholds stol=0.7, angle_tol=5, and ltol=0.2 is utilized. The RMSE is computed between the matched structure and ground truth, normalized by $3]{V/N}$, where $V$ and $N$ represent the lattice volume and atom numbers in the unit cell, respectively. The RMSE can also be obtained through the StructureMatcher class in pymatgen.
@@ 2.3 CPCP Module Aligns PXRD Space with Crystal Structure Space
To qualitatively assess the efficacy of the embeddings obtained from the CPCP module, we employ t-SNE to reduce the PXRD feature embeddings of hMOF-100 dataset into two dimensions, which is shown in Fig. 2a. The normalized volume of the unit cell corresponding to the PXRD crystal structure is represented in blue, with colors closer to blue indicating larger volumes and colors closer to white indicating smaller volumes. The clustering of colors demonstrates that similar PXRD embeddings share similar volumes. As the unit cell volume is a representative property of crystal structures, the PXRD embeddings effectively capture underlying structural characteristics, thereby aligning the PXRD space with the crystal structure space to a certain degree.
@@
More specifically, structure embeddings from the test sets are compiled by the crystal structure network to create a retrieval database. Searches are conducted using PXRD embeddings produced by the PXRD feature extractor as query input, employing cosine similarity as the search metric. The search results are summarized in Fig. 2b. The top-10 hit rate reaches 94.4% in the hMOF-100 dataset and 74.3% in the hMOF-400 dataset.
@@
To gain a more intuitive understanding of the retrieval effectiveness, we randomly select 50 crystal structures and their corresponding PXRD patterns, using their similarity scores to construct a heatmap as displayed in Fig. 2c. In the heatmap, colors closer to blue indicate higher similarity, while colors closer to white signify lower similarity. The heatmap’s diagonal represents the correctly matched pairs. A distinct blue distribution along the diagonal demonstrates the model’s ability to accurately identify crystal structures corresponding to the PXRD patterns.
@@
Additionally, we visualize a retrieval result for a given PXRD pattern. The four highest-scoring crystal structures retrieved by the CPCP module, their corresponding PXRD patterns, and their ranks are presented in Fig. 2d. It is evident that the non-ground-truth structures with top ranks exhibit similar metal-connecting structures, with the connected ligands also displaying a high degree of similarity. From the perspective of PXRD spectra, the peak positions of the high-intensity peaks in the PXRD spectra associated with similar structures also exhibit high similarity. Consequently, our retrieval model effectively extracts high-level information corresponding to the structures from the PXRD spectra, laying the foundation for the subsequent generation model.
2.4 XtalNet Can Generate Crystal Structure Conditioned on PXRD

### v3

@@ Abstract
Powder X-ray diffraction (PXRD) is a prevalent technique in materials characterization. While the analysis of PXRD often requires extensive human manual intervention, and most automated method only achieved at coarse-grained level. The more difficult and important task of fine-grained crystal structure prediction from PXRD remains unaddressed. This study introduces XtalNet, the first equivariant deep generative model for end-to-end crystal structure prediction from PXRD. Unlike previous crystal structure prediction methods that rely solely on composition, XtalNet leverages PXRD as an additional condition, eliminating ambiguity and enabling the generation of complex organic structures with up to 400 atoms in the unit cell. XtalNet comprises two modules: a Contrastive PXRD-Crystal Pretraining (CPCP) module that aligns PXRD space with crystal structure space, and a Conditional Crystal Structure Generation (CCSG) module that generates candidate crystal structures conditioned on PXRD patterns. Evaluation on two MOF datasets (hMOF-100 and hMOF-400) demonstrates XtalNet’s effectiveness. XtalNet achieves a top-10 Match Rate of 90.2% and 79% for hMOF-100 and hMOF-400 in conditional crystal structure prediction task, respectively. XtalNet enables the direct prediction of crystal structures from experimental measurements, eliminating the need for manual intervention and external databases. This opens up new possibilities for automated crystal structure determination and the accelerated discovery of novel materials.
keywords:
Deep Learning, Powder X-ray Diffraction, Crystal Structure Prediction, Equivariant Deep Generative Model, Metal-organic Frameworks (MOFs)
Qingsi Lai Fanjie Xu Lin Yao* Zhifeng Gao Siyuan Liu Hongshuai Wang Shuqi Lu Di He Liwei Wang Linfeng Zhang Cheng Wang* Guolin Ke*
[…]
{affiliations}
Q. Lai, F. Xu, L. Yao, Z. Gao, S. Liu, H. Wang, S. Lu, L. Zhang, G. Ke
DP Technology, Beijing, 100080, China
Email Address: yaol@dp.tech; kegl@dp.tech
Q. Lai, L. Wang
Center for Data Science, Peking University, Beijing 100871, China
D. He, L. Wang
School of Intelligence Science and Technology, Peking University, Beijing 100871, China
F. Xu, C. Wang
College of Chemistry and Chemical Engineering, Xiamen University, Xiamen, 361005, China
Email Address: wangchengxmu@xmu.edu.cn
L. Zhang, C. Wang
AI for Science Institute, Beijing, 100084, China
1 Introduction
Powder X-ray Diffraction (PXRD) [1] is widely used material characterization techniques for its cost-effectiveness and its ability to analyze crystallography [2]. PXRD produces a pattern that encodes information about the crystal symmetry, lattice parameters and atomic positions [3]. Comparing XRD patterns of candidate materials with the measured or simulated XRD patterns of known materials allows for deriving all the aforementioned information. However, the analysis of PXRD patterns often involves multiple sequential steps and extensive manual intervention, leading to several attempts to automate PXRD analysis [4, 5, 6, 7, 8, 9]. Although these efforts have achieved success at coarse-grained levels, such as phase identification [5] and space group prediction [9], the more challenging task of fine-grained crystal structure prediction from PXRD remains unaddressed. In this paper, we propose an end-to-end method named XtalNet to identify crystal structures from PXRD.
@@ 1 Introduction
In PXRD analysis, determining the crystal structure of experimental materials is one of the most important and challenging steps. Existing PXRD-based crystal structure determination methods involve comparing experimental PXRD patterns of unknown materials against databases like the Inorganic Crystal Structure Database (ICSD) [10] to identify structural analogues as a starting point. Subsequently, Rietveld refinement[11] is employed to incrementally refine the rough analogous structure and achieve greater precision. However, the incompleteness of the database and the complexity of the procedures complicate the structure determination process. Additionally, the Rietveld refinement step often requires extensive manual intervention from experienced scientists. Therefore, a method that can directly predict crystal structures from PXRD data without relying on external databases has the potential to significantly reduce the manual labor required by chemists and represent a major stride towards automating PXRD analysis.
Crystal structure plays a critical role in various scientific realms such as physics, chemistry, and material science, as many physical and chemical properties of materials are determined by the crystal structure [12, 13, 14]. Crystal structure prediction (CSP), which aims to obtain the three-dimensional structure of crystals based on their composition [15], has made significant progress with machine learning [16, 17, 18, 19, 20, 21, 22, 23]. Most of these methods are carefully designed to account for the unique challenges of periodicity and equivariance in CSP, and some of them [20, 21, 22] utilize diffusion models [24, 25] to achieve remarkable structure generation results. However, these methods have been primarily tested on inorganic crystal structure datasets such as Perov-5 [26, 27], MP-20 [28], and ICSD [29], which consist of a limited number of atoms in the unit cell (usually less than 50). The efficacy of these methods in predicting more complex and larger organic structures or under stringent conditions has yet to be validated. Furthermore, the CSP problem is typically formulated to predict the minimum energy structure. However, as the number of atoms in a unit cell increases, one composition may correspond to many crystal structures as the energy landscape becoming complex. Therefore, a more practical scenario is predicting a material’s crystal structure based on its experimental characterization. The existence of this characterization leads to a one-to-one correspondence with crystal structure. Thus, integrating experimental characterization into crystal structure prediction is a more practical application scenario. Previous crystal structure prediction methods provide us with a powerful tool for generating crystal structures. Based on this powerful tool, what we need to consider is how to use experimental characterization as a condition to guide the crystal structure prediction.
While PXRD is fundamentally different from crystal structure, a sole prediction module is not able to achieve finer structure prediction that satisfies the PXRD condition, as shown in Results 2.4.1. A prediction module is required to predict the crystal structure and align the PXRD latent space with the crystal space simultaneously, which poses a significant challenge. Therefore, incorporating prior information into the prediction module to provide alignment between the PXRD space and crystal structure space is necessary. A common method is to use pretraining, which obtains a post-aligned PXRD encoder as the initialization of the prediction module. The necessity of the pretraining procedure for multi-modality tasks has been validated in many other fields, such as text-to-image generation [30, 31, 32, 33], with prominent examples such as Stable Diffusion [34]. These models employ text embeddings generated by pretrained CLIP [35] text encoders to guide the diffusion process in generating images conditioned on textual descriptions. In the case of crystal structure prediction, PXRD serves as the key signal to guide the generation of crystal structures. Drawing inspiration from these successful multi-modality models in other domains, we have explored the potential of integrating a contrastive learning approach into our conditional crystal structure prediction methodology, resulting in the contrastive PXRD-crystal pre-training (CPCP) module. By pretraining the PXRD encoder using the Contrastive PXRD-Crystal Pre-training (CPCP) module, PXRD signals closer to the crystal structure space can be provided. The PXRD encoder pretraining by the CPCP module can guide the generation of crystal structures.
Therefore, we propose XtalNet, the first equivariant deep generative model for end-to-end crystal structure prediction from PXRD. XtalNet aims to extend the capabilities of deep learning in predicting crystal structures based on PXRD patterns, encompassing more complex structures and specific conditions. Unlike previous CSP methods that solely rely on composition for obtaining crystal structures, XtalNet incorporates PXRD as a supplemental condition, ensuring a one-to-one mapping without ambiguity, which can be applied in a real experimental setting. To the best of our knowledge, XtalNet is the first deep learning method to directly predict the fine-grained crystal structure, which makes it different from previous PXRD analysis methods that only predict coarse-grained properties. XtalNet is an end-to-end deep learning based method that eliminates the need for human intervention or external database dependencies, distinguishing it from traditional PXRD analysis approaches involving complex multi-step procedures and human insights. Our model is capable of handling organic crystal systems, such as metal-organic frameworks (MOFs) that are important for gas separation [36], even when the unit cell contains a large number of atoms (up to 400). To achieve this, we have compiled PXRD-crystal datasets named hMOF-100 and hMOF-400 with more than 100,000 data points.
Our approach is made possible through the contrastive PXRD-crystal pre-training (CPCP) module and the conditional crystal structure generation (CCSG) module. The CPCP module employs contrastive learning pre-training to align the PXRD space with the crystal structure space. Furthermore, by utilizing the CCSG modules, multiple candidate crystal structures can be generated conditioned on the PXRD pattern by equivariant diffusion model. These candidate structures are subsequently scored and ranked using the CPCP module, thus accomplishing the ranked structure prediction task with a top-10 match rate of 90.2% in the hMOF-100 dataset and 79% in the hMOF-400 dataset.
@@ 2.1 Overview of XtalNet
XtalNet is designed to predict crystal structures from PXRD pattern, which we approach as a conditional generation task. The goal is to generate the corresponding crystal structure based on the given PXRD pattern. To achieve this, we have developed two key modules, as depicted in Figure 1a and b: the Contrastive PXRD-Crystal Pretraining (CPCP) module and the Conditional Crystal Structure Generation (CCSG) module. The CPCP module primarily aligns the crystal space with the PXRD space, and the pre-trained PXRD feature extractor is subsequently used to initialize the CCSG module. The CCSG module is specifically designed to reconstruct the crystal structure from a given PXRD pattern using the pre-trained PXRD feature extractor derived from the CPCP module by using equivariant diffusion model.
Figure 1: Overview of XtalNet. a, Framework of the Contrastive PXRD-Crystal Pretraining (CPCP) module. The CPCP module takes PXRD patterns and crystal structures as inputs and produces similarity scores between them. A transformer-based PXRD feature extractor processes the PXRD pattern data, while an equivariant Graph Neural Network (GNN) extracts features from the crystal structure. The similarity score is computed as the dot product of these two feature sets. b, Framework of the Conditional Crystal Structure Generation (CCSG) module. The CCSG module utilizes the PXRD pattern as a condition to generate crystal structures. The PXRD feature extractor is initialized from the CPCP module pretraining and kept frozen. Subsequently, the composition of the crystal are used to initialize the atom positions and lattice matrix. The denoising network, referred to as the crystal structure network, takes the previous crystal structure, along with the PXRD feature obtaind through PXRD feature extractor and time step, as inputs to update the crystal structure. This process is iteratively repeated in a reverse manner. c, PXRD data generation workflow. PXRD data can be acquired in two ways: by simulating PXRD data from a given crystal structure using GSAS [37] software, or by conducting actual PXRD experiments with an XRD instrument. d, Workflow of the PXRD feature extractor. The PXRD data is first tokenized into peak tokens based on peak intensity and the corresponding diffraction angle, after which PXRD features are derived using BERT. e, Framework of the crystal structure network. In the CPCP model, only the solid line components of the process are executed, whereas in the CCSG model, both solid and dashed line components are executed.
As shown in Figure 1a, the CPCP module is inspired by the design of the CLIP model [35]. It employs a transformer-based [38] PXRD feature extractor to encode PXRD patterns into feature representations. A crystal structure feature network, based on Equivariant Graph Neural Networks [21], is then utilized to extract crystal structure features. The similarity between PXRD features and crystal structure features is calculated using cosine similarity. Matching pairs of PXRD patterns and corresponding crystal structures are treated as positive pairs, while non-matching pairs are considered negative. Contrastive learning is performed using the InfoNCE loss function. The framework of CCSG module is shown in Figure 1b. The CCSG module operates within a diffusion-based framework, where the PXRD feature serves as a key condition for generation. The reverse processes is used for inference, which aims to update the noised crystal structure. Hence the noised crystal structure, PXRD feature and time embedding are all passed through the crystal structure network to denoise the fractional coordinates and lattice. The diffusion process enables the CCSG training. Noise is added to the crystal structure via the diffusion process, and the goal of the crystal structure network is to predict this noise. It is worth mentioning that the PXRD feature extractor is initialized by the pretraining of CPCP module, which is vital for generation performance. While multiple crystal structure candidates can be sampled from CCSG module, the CPCP module can be used to rank them by the similarity between PXRD pattern and candidates. This process enables us to screen multiple candidates effectively.
The architecture of the PXRD feature extractor is illustrated in Figure 1d. Initially, the PXRD pattern is tokenized based on its peaks, and a [CLS] token, representing the global PXRD feature, is added at the header of perak tokens. A BERT model [39] is used to obtain the PXRD features, with the global feature from the header being utilized in both the CPCP and CCSG modules. The crystal structure network is employed in both modules, though with some differences, as shown in Figure 1e. In the CPCP module, only the composition is used as the node feature. However, in the CCSG module, the PXRD feature and time embedding are also used as node features, concatenated with the composition embedding. The output of the crystal structure network in the CPCP module is the final graph feature, used for contrastive learning, while in the CCSG module, the output includes both the coordinate score and lattice noise. The coordinate score is obtained from the final node feature after MPNN processing, and the lattice noise is derived from the final graph feature and the original lattice. The crystal structure networks in the two modules do not share parameters and are each trained independently from scratch. More details are discussed in Method 4.
@@
To accomplishing XtalNet training, we must obtain pairs of crystal structure and PXRD. The entire data workflow is illustrated in Figure 1c. While crystal structure is prevalent in various databases, PXRD patterns are not ordinarily included in standard databases. Since high-quality experimental data is extremely rare and costly, we simulate PXRD patterns from known crystal structures using GSAS [37] software as our training data, which can generate a large amount of data at limited cost. At the same time, we also use X-ray diffraction instruments to collect a limited number of experimental PXRD patterns to test XtalNet performance in practice as a case study.
@@ 2.2 Dataset Preparing and Evaluation Metrics
In order to evaluate the CPCP module, the database retrieval task is designed, which aims to identify matching crystal structures from a given set of candidate structures based on known PXRD pattern. As the purpose of CPCP module is to obtain matching relationship between PXRD space and crystal structure space, the database retrieval task is suitable for validating it. Hence, We employ the top-$k$ hit rate as the evaluation metric, which measures the frequency at which the desired crystal structures are discovered. For the evaluation of CCSG module, following previous works [20, 21], we assess the crystal structure using Match Rate and RMSE. Specifically, the Match Rate denotes the proportion of matched generated structures relative to all ground truth. The StructureMatcher class in pymatgen [45] is utilized. The RMSE is computed between the matched structure and ground truth, normalized by $3]{V/N}$, where $V$ and $N$ represent the lattice volume and atom numbers in the unit cell, respectively. The RMSE can also be obtained through the StructureMatcher class in pymatgen.
@@ 2.3 CPCP Module Aligns PXRD Space with Crystal Structure Space
To qualitatively assess the efficacy of the embeddings obtained from the CPCP module, we employ t-SNE to reduce the PXRD feature embeddings of hMOF-100 dataset into two dimensions, which is shown in Figure 2a. The normalized volume of the unit cell corresponding to the PXRD crystal structure is represented in blue, with colors closer to blue indicating larger volumes and colors closer to white indicating smaller volumes. The clustering of colors demonstrates that similar PXRD embeddings share similar volumes. As the unit cell volume is a representative property of crystal structures, the PXRD embeddings effectively capture underlying structural characteristics, thereby aligning the PXRD space with the crystal structure space to a certain degree.
@@
More specifically, structure embeddings from the test sets are compiled by the crystal structure network to create a retrieval database. Searches are conducted using PXRD embeddings produced by the PXRD feature extractor as query input, employing cosine similarity as the search metric. All experiments are conduct on the whole test set of hMOF-100 and hMOF-400 dataset. The search results are summarized in Figure 2b. The top-10 hit rate reaches 97.2% in the hMOF-100 dataset and 74.1% in the hMOF-400 dataset. Here is a performance decline from hMOF-400 to hMOF-100, which can be attributed to two main reasons. First, the model faces a more complex system, resulting in a larger search space and an increased difficulty in making accurate predictions, leading to reasonable poorer performance. Second, there is relatively less data available for larger systems compared to smaller ones, and the model has not been sufficiently trained on the large system data, which also contributes to worse results.
@@
To gain a more intuitive understanding of the retrieval effectiveness, we randomly select 50 crystal structures and their corresponding PXRD patterns, using their similarity scores to construct a heatmap as displayed in Figure 2c. In the heatmap, colors closer to blue indicate higher similarity, while colors closer to white signify lower similarity. The heatmap’s diagonal represents the correctly matched pairs. A distinct blue distribution along the diagonal demonstrates the model’s ability to accurately identify crystal structures corresponding to the PXRD patterns.
@@
Additionally, we visualize a retrieval result for a given PXRD pattern. The four highest-scoring crystal structures retrieved by the CPCP module, their corresponding PXRD patterns, and their ranks are presented in Figure 2d. It is evident that the non-ground-truth structures with top ranks exhibit similar metal-connecting structures, with the connected ligands also displaying a high degree of similarity. From the perspective of PXRD spectra, the peak positions of the high-intensity peaks in the PXRD spectra associated with similar structures also exhibit high similarity. Consequently, our retrieval model effectively extracts high-level information corresponding to the structures from the PXRD spectra, laying the foundation for the subsequent generation model.
2.4 XtalNet Can Predict Crystal Structure Conditioned on PXRD